OpenStudy (anonymous):

find the product 7x(-2x^4)

OpenStudy (anonymous):

Could you show us what you've done so far?

OpenStudy (anonymous):

i just cant seem to understand

OpenStudy (anonymous):

Okay, well your above expression can be rewritten as this: \[\large 7x(-2x^4)=7\times x\times -2\times x^4\] AM I right?

OpenStudy (anonymous):

Or do you not understand this part?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I need to find the product

OpenStudy (anonymous):

This is what we're doing. I'm getting there step by step for you to understand. So basically what the expression is asking you to do is calculate \[\huge 7×x×−2×x^4=(7\times -2)\times (x\times x^4)\]

OpenStudy (anonymous):

So could you simplify the RHS [Right Hand Side] for me?

OpenStudy (anonymous):

what's 7 times -2? And what's x\times x^4. Use this identity for the exponents: \[\huge x^n\times x=x^{n+1}\] where: n is a real number.

OpenStudy (anonymous):

@amandahigham I'm waiting for your response to what I put out for you.